OpenStudy (anonymous):

graph and show calculations 3x+5y ≤ 15

OpenStudy (anonymous):

here is my graph is this correct? not sure what to put for calculation

OpenStudy (cwrw238):

the slope is - 3/5

OpenStudy (anonymous):

how did you find the slope for that?

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

if you rearrange 5y = -3x + 15 y = (-3/5) x + 3

OpenStudy (cwrw238):

your slope = -4.5/7.5 = -3/5 -correct

OpenStudy (cwrw238):

3x + 5y = 15 when x = 0 y = 3 when y = 0 x = 5

OpenStudy (cwrw238):

yes that checks out and you have a continuous line not a broken one looks good